OpenStudy (anonymous):

You start out with $20 and then spend money in a store where every item is $3. Y = -3x + 20. Use your graph for problem to predict the answer. What is the largest number of items you can buy?

OpenStudy (anonymous):

Notice that y = -3(7) + 20 or Y = -21 + 20 = -1 Can we have negative money here? No credit cards!! If we cannot have negative money, 7 does not work, so the most we can buy is ?

OpenStudy (anonymous):

41?

OpenStudy (anonymous):

almost .... if you cannot afford 7 items, the store will not sale you 41 items. If you cannot afford 7 items, then you can buy? how many?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

6. If you cannot afford 7, then you can buy one less, 6.

OpenStudy (anonymous):

lol oh well thank you for telling me

OpenStudy (anonymous):

your welcome